# The Analytic Stockwell Transform and its Zeros

### v2

@@ Abstract
A recent original line of research in time–frequency analysis has shifted the interest in energy maxima toward zeros. Initially motivated by the intriguing uniform spread of the zeros of the spectrogram of white noise, it has led to fruitful theoretical developments combining probability theory, complex analysis and signal processing. In this vein, the present work proposes a characterization of the zeros of the Stockwell Transform of white noise, which consists in an hybrid time–frequency multiresolution representation. First of all, an analytic version of the Stockwell Transform is designed. Then, analyticity is leveraged to establish a connection with the hyperbolic Gaussian Analytic Function, whose zero set is invariant under the isometries of the Poincaré disk. Finally, the theoretical spatial statistics of the zeros of the hyperbolic Gaussian Analytic Function and the empirical statistics of the zeros the Analytic Stockwell Transform of white noise are compared through intensive Monte Carlo simulations, supporting the established connection. A publicly available documented Python toolbox accompanies this work.
@@ 1 Introduction
Time–frequency analysis is a cornerstone of nonstationary signal processing [13, 18, 16]. Crucially, contrary to the Fourier spectrum describing the overall frequency content but loosing time localization, time–frequency representations aim at accounting for both the temporal dynamics and the time-varying frequency content of a signal. Among the most widely used time–frequency representations are the spectrogram, obtained as the squared modulus of the Short-Time Fourier Transform [13, 16], and the time–scale scalogram, which is the squared modulus of the Wavelet Transform [34]. These joint representations appeared particularly suited to analyze physical phenomena as diverse as bat echolocation [14], earthquakes [42] and gravitational waves [8].
@@
The search for analytic signal transforms is a long-standing problem both in signal processing and harmonic analysis [4, 26, 39, 20, 6, 30] as analyticity is a key property in phase retrieval problems and phaseless reconstruction [26, 39, 20]. Recently, this interest has been strongly renewed by an unorthodox research path which emerged at the interface between time–frequency analysis and stochastic geometry. This novel line of research consists in shifting the historical interest in representation maxima toward zeros, which can be seen as “silent" points in the time–frequency plane [17]. Notably, the seminal works [17, 15] have shed light on the remarkably even distribution of the zeros of the spectrogram white noise. Not only this observation received a rigorous mathematical explanation [5] elegantly combining complex analysis and probability theory, but the established link between time–frequency analysis and random complex analytic functions opened up a novel line of research in the theory of signal representations and led to the characterization of the zeros of many other signal representations, among which the Wavelet Transform [3, 24, 5, 6]. Further, this fruitful connection has even motivated the design of new representations, and in particular of discrete representations with desirable algebraic properties [6, 30], as well as original contributions in harmonic analysis [2].
@@
The present work aims at completing these studies by characterizing the zeros of the time–frequency Stockwell Transform [37], which genuinely hybridizes the Short-Time Fourier Transform and the Wavelet Transform so as to provide an original multi-resolution analysis with a local phase having an absolute reference [37, 36, 33]. The motivation for the introduction of the Stockwell transform was to overcome a major practical drawback of the Short-Time Fourier Transform: for low (resp. high) frequency the frequency (resp. time) resolution is coarse, impairing precise localization in the time-frequency plane. To that aim the historical design relies on a Gaussian analysis window [37], whose width is varied as a function of frequency so as to produce a higher frequency resolution at lower frequencies and a higher time resolution at higher frequencies while preserving the absolute reference of the phase, which is lost when using the Wavelet Transform. Several extensions of the Stockwell Transform have been proposed in the literature [32, 7, 25, 29, 43], to cite but a few. The main objective of these extensions is to enable to adapt closely to the characteristics of the analyzed signal and to the time–frequency patterns of interest.
Thanks to its solid theoretical ground, high versatility and ease of interpretation, the Stockwell Transform has progressively gained popularity not only in signal processing [40, 22, 41], but also in applied mathematics [33, 35] and mathematical physics [19, 23].
@@
The main contributions of the present work are, first, to construct an Analytic Stockwell Transform; second, thanks to a rigorous construction of an adapted continuous white noise, to characterize the zeros of the Analytic Stockwell Transform of white noise by proving that they coincides with the zeros of the so-called hyperbolic Gaussian Analytic Function hence demonstrating that the zeros are invariant under isometries of the Poincaré disk; third, to provide numerical evidence supporting this connection by comparing through intensive Monte Carlo simulations the theoretical pair correlation function of the zeros of the hyperbolic Gaussian Analytic Function and the empirical pair correlation function of the zeros of the Analytic Stockwell Transform of white noise. In Section 2, after reminding the definition of the Generalized Stockwell Transform, the existence of an Analytic Stockwell Transform is proven by leveraging the intrinsic link between the Stockwell and Wavelet transforms. Then, the distribution of the zeros of the Analytic Stockwell Transform of white noise is characterized in Section 3, taking advantage of a well-suited construction of white noise and of a connection with Gaussian Analytic Functions. Finally, numerical experiments are implemented in Section 4, providing illustrations of the zero pattern of the Analytic Stockwell Transform and quantitative support to the results of Section 3 by systematically comparing the theoretical and empirical spatial statistics of zeros in hyperbolic geometry through intensive Monte Carlo simulations.
@@ 2 The Analytic Stockwell Transform
First, the formal definition of the Generalized Stockwell Transform is reminded, as well as its interpretation as a covariant transform under the group generated by time translations, frequency modulations and scale dilatations. Then, connecting the Stockwell Transform to the Wavelet Transform and leveraging analyticity results obtained for Cauchy wavelet, an Analytic Stockwell Transform is constructed.
@@ 2.1 Preliminary: definition of the Generalized Stockwell Transform
Let $\varphi\in L^{2}(\mathbb{R})$ a finite energy window. The Generalized Stockwell Transform with analysis window $\varphi$ of a finite energy signal, $f\in L^{2}(\mathbb{R})$, is defined as
@@
Then, the Generalized Stockwell Transform of Equation (1) can be interpreted as the decomposition of the signal onto the covariant family obtained by making the group generated by translations, modulations and dilatations acting on the analysis window, that is:
@@ 2.2 Existence of an Analytic Stockwell Transform
The design of an Analytic Stockwell Transform leverages the Cauchy wavelet, also referred to as the Daubechies-Paul wavelet in the literature [11].
@@
Let $\beta\in\mathbb{R}$ be such that $\beta>0$. The Cauchy wavelet of parameters $\beta$ is defined as the unique function $\psi_{\beta}\in L^{2}(\mathbb{R})$ whose Fourier transform writes
@@
Let $\beta>0$ and $\psi_{\beta}\in L^{2}(\mathbb{R})$ be the Cauchy wavelet of parameter $\beta$. Define the modulated Cauchy wavelet $\varphi_{\beta}(t)=\psi_{\beta}(t)\mathrm{e}^{-2\pi\mathrm{i}t}$. For any finite energy signal $f\in L^{2}(\mathbb{R})$, there exists an analytic function on the unit disk $F:\mathbb{D}\rightarrow\mathbb{C}$, a conformal mapping $\vartheta:\mathbb{C}_{+}\rightarrow\mathbb{D}$, and a smooth nonvanishing function $\lambda:\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{C}$ such that
$\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\quad S_{\varphi_{\beta}}f(x,\xi)=\lambda(x,\xi)\times F\left(\vartheta(x+\mathrm{i}\xi^{-1})\right).$ (8)
@@
Then, from [26, Section 2.2] and  [20, Theorem 2], the Cauchy wavelet transform of $f\in L^{2}(\mathbb{R})$ is analytic on the upper-half complex plane, that is $G:z=x+\mathrm{i}y\mapsto y^{-\beta-1/2}W_{\psi_{\beta}}f(x,y)$ is an analytic function on $\mathbb{C}_{+}$. Moreover, let $\vartheta:\mathbb{C}_{+}\rightarrow\mathbb{D}$ be the Cayley transform defined as
@@
Define $F:\mathbb{D}\rightarrow\mathbb{C}$ as $\forall w\in\mathbb{D},\,F(w)=G(\vartheta^{-1}(w))$. $F$ being the composition of an analytic function $G$, and of a conformal, hence holomorphic, mapping $\vartheta^{-1}$, $F$ is analytic. Finally, let $\lambda(x,\xi)=\xi^{-\beta}\mathrm{e}^{-2\pi\mathrm{i}\xi x}$, then $\lambda:\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{C}$ is a smooth nonvanishing function.
By construction of $F$, $\vartheta$ and $\lambda$, they satisfy all the conditions enumerated in Theorem 1 and $\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\,S_{\varphi_{\beta}}f(x,\xi)=\lambda(x,\xi)\times F\left(\vartheta(x+\mathrm{i}\xi^{-1})\right)$ which concludes the proof. ∎
@@ 3 The Zeros of the Analytic Stockwell Transform of White Noise
For the sake of completeness, before proving the connection with the hyperbolic Gaussian Analytic Function, some necessary elements of hyperbolic and stochastic geometry are presented; the definition of the hyperbolic Gaussian Analytic Function is reminded accompanied with its main interesting properties so that the present section is self-content. The interested reader can refer to [21] for a thorough presentation of the theory of Gaussian Analytic Functions and a detailed study of their zeros.
@@ 3.1 Gaussian Analytic Functions and point processes
The Poincaré disk model is the triplet $(\mathbb{D},\mathsf{d}_{\mathbb{D}},\mathsf{m}_{\mathbb{D}})$ constituted of the open unit disk $\mathbb{D}=\{z\in\mathbb{C}:\lvert z\rvert<1\}$, the hyperbolic metric defined as
$\mathrm{d}\mathsf{m}_{\mathbb{D}}(z)=\frac{\mathrm{d}z}{\left(1-\lvert z\rvert^{2}\right)^{2}}$ (12)
where $\mathrm{d}z$ refers to the standard Lebesgue measure on $\mathbb{C}$, and the associated hyperbolic distance defined for $z,w\in\mathbb{D}$ as
@@
Let $\alpha>0$ a real number, the hyperbolic Gaussian Analytic Function is defined on $\mathbb{D}$ as
@@
where $\left(\zeta_{n}\right)_{n\in\mathbb{N}}$ is a sequence of independent identically distribution standard Gaussian random variables, $\zeta_{n}\sim\mathcal{N}_{\mathbb{C}}(0,1)$. By [21, Lemma 2.2.3], for $\alpha>0$, $\mathsf{GAF}_{\mathbb{D}}^{(\alpha)}$ is well-defined and almost-surely an analytic function on $\mathbb{D}$.
Gaussian Analytic Functions have received much interest from the probability and spatial statistics communities, in particular due to the remarkable properties of their zeros [21]. Indeed, since Gaussian Analytic Functions are almost-surely analytic, their zeros constitute a random set of isolated points in $\mathbb{C}$ with probability one. Further, [21, Lemma 2.4.1] ensures that the random zeros of a Gaussian Analytic Function are almost-surely simple. Hence, the zero set of a Gaussian Analytic Function forms a simple point process [10, 28], defined as a random variable taking values in the configurations of points in $\mathbb{C}$. Several Gaussian Analytic Functions, among which the hyperbolic Gaussian Analytic Function, are distinguished by the fact that their zeros are distributed very uniformly in their definition domain, which is mathematically formalized as an invariance under isometries [21, Section 2.3]. This invariance property is the cornerstone of the zero-based signal processing procedures designed for detection [5, 30], denoising [24, 5] and component separation [15]. The interested reader can refer to [31] for a review of the use of spatial statistics in time–frequency analysis.
@@
[21, Proposition 2.3.4] For $\alpha>0$, the point process formed by the zeros of $\mathsf{GAF}_{\mathbb{D}}^{(\alpha)}$ is invariant under the isometries of the hyperbolic disk.
@@ 3.2 Link with the zeros of the Hyperbolic Gaussian Analytic Function
A major consequence of Theorem 1 is that, since $\vartheta$ is invertible and $\lambda$ never vanishes, the zeros of the Analytic Stockwell Transform of a finite energy signal are in one-to-one correspondance with the zeros of an analytic function $F$ on the unit disk. As so, the zeros of the Analytic Stockwell Transform form a collection of isolated points in $\mathbb{D}$. Further, in the presence of noise in the signal, $f$ is random and then the zeros of $S_{\varphi_{\beta}}f$ are random isolated points in $\mathbb{D}$, thus form a point process.
In line with [3, 6, 30], the purpose of this section is to study a remarkable instance of such point process, that is the zeros of the Analytic Stockwell Transform of white noise, and to demonstrate that it coincides in law with the zeros of the hyperbolic Gaussian Analytic Function introduced in Section 3.1. Following [3, 24, 5, 6], this connection is established in two steps: first the continuous white noise adapted to the considered transform is rigorously constructed; second the equality, in law, of the point process of the zeros of the transform of white noise and of the zeros of the associated Gaussian Analytic Function is proven. The following lemma will appear as an important building block for the design of the continuous white noise adapted to the Analytic Stockwell Transform.
@@
There exists a smooth nonvanishing function $\eta_{\beta}:\mathbb{R}\times\mathbb{R}_{+}$ such that, for all $n\in\mathbb{N}$, the Analytic Stockwell Transform of $f_{n}$ has the closed-form expression:
@@
The above lemma is an original contribution of the present work. Indeed, first, the Hilbertian basis used to connect the Cauchy Wavelet Transform of white noise to the Hyperbolic Gaussian Analytic Function in [3, 24, 6] has to be adapted to the Analytic Stockwell Transform of Section 2.2. Second, in [3, 24] the Hilbertian basis used is only defined implicitly. Third, it seems that the explicit derivations performed in [6, Theorem 2.3], relying on [6, Equation (2.15)], together with the computations in [1, Section 4], contains erroneous prefactors that might impact the definition of the white noise adapted to the Cauchy wavelet transform used to prove [6, Theorem 2.3]. Hence, since Equation (18) is the cornerstone of the demonstration of the connection between the hyperbolic Gaussian Analytic Function and the Analytic Stockwell Transform of white noise, for the sake of completeness, the proof below details the entire calculation, including the proposed corrected prefactors and normalizations.
@@
involves the convolution between $f$ and $h$ defined as $h(t)=\overline{\psi_{\beta}(-t/\xi^{-1})}$. Then, using that the Fourier transform of the convolution between two functions is the product of the Fourier transforms of these functions leads to
@@
which is the Laplace transform [^1] Let $g:\mathbb{R}\rightarrow\mathbb{C}$. The Laplace transform of $g$, denoted $\mathcal{L}[g]:\mu\subseteq\mathbb{C}\rightarrow\mathbb{C}$, is a function of the complex variable $s$ defined as: $\mathcal{L}\{g\}(s)=\int_{\mathbb{R}}\mathrm{e}^{-\nu s}g(\nu)\,\mathrm{d}\nu$, $s\in\mu$. of the $n$th order derivative of $g_{n}(\nu)=\mathrm{e}^{-\nu}\nu^{2\beta+n}$ evaluated in $-(\mathrm{i}z+1)/2$, i.e.,
@@
$\begin{split}\mathcal{L}[g_{n}](s)=\int_{0}^{\infty}\mathrm{e}^{-s\nu}\mathrm{e}^{-\nu}\nu^{2\beta+n}\,\mathrm{d}\nu&=\frac{\Gamma(2\beta+n+1)}{(1+s)^{2\beta+n+1}}\end{split}$ (25)
leveraging the generalized change of variable $\nu^{\prime}=\nu(1+s)$ and the definition of the Euler Gamma function. The second step is to compute the limits $g^{(k)}(0^{+})$ for $k\in\{0,\ldots,n-1\}$. To that aim, a recursion shows that for all $k\in\{0,1,\ldots,n\}$, there exists a polynomial $p_{k}$ of order $n$, such that
@@
where $\vartheta$ is the Cayley transform defined in Equation (10). Finally, remembering that $z=x+\mathrm{i}\xi^{-1}$, let $\eta_{\beta}:\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{C}$ be defined as
$\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\quad\eta_{\beta}(x,\xi)=\sqrt{\xi}\mathrm{e}^{-2\pi\mathrm{i}\xi x}\left(\frac{\mathrm{i}}{x+\mathrm{i}\xi^{-1}+\mathrm{i}}\right)^{2\beta+1}$ (29)
Then, for any $\beta>0$, $\eta_{\beta}$ is a well-defined a smooth nonvanishing function on the upper half-plane, and the Analytic Stockwell Transform of $f_{n}$ writes

### v3

@@ Abstract
The Stockwell Transform is a time–frequency representation resulting from an hybridization between the Short-Time Fourier Transform and the Continuous Wavelet Transform. Instead of focusing on energy maxima, an unorthodox line of research has recently shed the light on the zeros of time–frequency transforms, leading to fruitful theoretical developments combining probability theory, complex analysis and signal processing. While the distributions of zeros of the Short-Time Fourier Transform and of the Continuous Wavelet Transform of white noise have been precisely characterized, that of the Stockwell Transform of white noise zeros remains unexplored. To fill this gap, the present work proposes a characterization of the distribution of zeros of the Stockwell Transform of white noise taking advantage of a novel generalized Analytic Stockwell Transform. First of all, an analytic version of the Stockwell Transform is designed. Then, analyticity is leveraged to establish a connection with the hyperbolic Gaussian analytic function, whose zero set is invariant under the isometries of the Poincaré disk. Finally, the theoretical spatial statistics of the zeros of the hyperbolic Gaussian analytic function and the empirical statistics of the zeros the Analytic Stockwell Transform of white noise are compared through intensive Monte Carlo simulations, showing the practical relevance of the established connection. A documented Python toolbox has been made publicly available by the authors.
@@ 1 Introduction
Time–frequency analysis is a cornerstone of nonstationary signal processing [14, 19, 17]. Crucially, contrary to the Fourier spectrum describing the overall frequency content but loosing time localization, time–frequency representations aim at accounting for both the temporal dynamics and the time-varying frequency content of a signal. Among the most widely used time–frequency representations are the spectrogram, obtained as the squared modulus of the Short-Time Fourier Transform [14, 17], and the time–scale scalogram, which is the squared modulus of the Wavelet Transform [35]. These joint representations appeared particularly well-suited to analyze physical phenomena as diverse as bat echolocation [15], earthquakes [44] and gravitational waves [9].
@@
The search for analytic signal transforms is a long-standing problem both in signal processing and harmonic analysis [4, 27, 41, 21, 7, 31] as analyticity is a key property in phase retrieval problems and phaseless reconstruction [27, 41, 21]. Recently, this interest has been strongly renewed by an unorthodox research path which emerged at the interface between time–frequency analysis and stochastic geometry. This novel line of research consists in shifting the historical interest in energy maxima toward zeros, which can be seen as “silent” points in the time–frequency plane [18]. Notably, the seminal works [18, 16] have shed light on the remarkably even distribution of the zeros of the spectrogram of white noise. Not only this observation received a rigorous mathematical explanation [6] elegantly combining complex analysis and probability theory, but the established link between time–frequency analysis and random complex analytic functions opened up a novel line of research in the theory of signal representation and led to the characterization of the zeros of many other signal representations, among which the Wavelet Transform [3, 25, 6, 7]. Further, this fruitful connection has even motivated the design of new representations, and in particular of discrete representations with desirable algebraic properties [7, 31], as well as original contributions in harmonic analysis [2].
@@
The present work aims at complementing these studies by characterizing the distribution of the zeros of the time–frequency Stockwell Transform [39], which genuinely hybridizes the Short-Time Fourier Transform and the Wavelet Transform so as to provide an original multi-resolution analysis with a local phase having an absolute reference [39, 38, 34]. The motivation for the introduction of the Stockwell transform was to overcome a major practical drawback of the Short-Time Fourier Transform: for low (resp. high) frequency the frequency (resp. time) resolution is coarse, impairing precise localization in the time-frequency plane. To that aim the historical design relies on a Gaussian analysis window [39] [37, Section 3.3], whose width is varied as a function of frequency so as to produce a higher frequency resolution at lower frequencies and a higher time resolution at higher frequencies while preserving the absolute reference of the phase, which is lost when using the Wavelet Transform. For a comprehensive and pedagogical presentation of the Wavelet and Stockwell Transforms, and links between both, the reader is referred to [35] and to the more recent textbook [37]. Several extensions of the Stockwell Transform have been proposed in the literature [33, 8, 26, 30, 45], to cite but a few. The main objective of these extensions is to enable to adapt closely to the characteristics of the analyzed signal and to the time–frequency patterns of interest. Thanks to its solid theoretical ground, high versatility and ease of interpretation, the Stockwell Transform has progressively gained popularity not only in signal processing [42, 23, 43], but also in applied mathematics [34, 36] and mathematical physics [20, 24].
@@
The main contributions of the present work are, first, to construct an Analytic Stockwell Transform; second, thanks to a rigorous construction of an adapted continuous white noise, to characterize the zeros of the Analytic Stockwell Transform of white noise by proving that they coincide with the zeros of the so-called hyperbolic Gaussian analytic function hence demonstrating that the zeros are invariant under isometries of the Poincaré disk; third, to provide numerical evidence supporting this connection by comparing through intensive Monte Carlo simulations the theoretical pair correlation function of the zeros of the hyperbolic Gaussian analytic function and the empirical pair correlation function of the zeros of the Analytic Stockwell Transform of white noise. In Section 2, after reminding the definition of the Generalized Stockwell Transform, the existence of an Analytic Stockwell Transform is proven by leveraging the intrinsic link between the Stockwell and Wavelet transforms. Then, the distribution of the zeros of the Analytic Stockwell Transform of white noise is characterized in Section 3, taking advantage of a well-suited construction of white noise and of a connection with Gaussian analytic functions. Finally, numerical experiments are implemented in Section 4, providing illustrations of the zero pattern of the Analytic Stockwell Transform and quantitative support to the results of Section 3 by systematically comparing the theoretical and empirical spatial statistics of zeros in hyperbolic geometry through intensive Monte Carlo simulations.
@@ 2 The Analytic Stockwell Transform
First, the formal definition of the Generalized Stockwell Transform is reminded. Then, connecting the Stockwell Transform to the Wavelet Transform and leveraging analyticity results obtained for Cauchy wavelet, an Analytic Stockwell Transform is constructed.
@@ 2.1 Preliminary: definition of the Generalized Stockwell Transform
Let $\varphi\in L^{2}(\mathbb{R})$. The Generalized Stockwell Transform with analysis window $\varphi$ of $f\in L^{2}(\mathbb{R})$ is defined as
@@
Then, the Generalized Stockwell Transform of Equation (1) can be interpreted as the decomposition of the signal onto family of functions generated by translations, modulations and dilatations of the analysis window, that is:
@@
Note the modulation operator $\mathbf{M}_{\xi}$, which essentially differentiates the Stockwell transform from the Wavelet transform and provides an absolute reference for the phase information [39].
@@ 2.2 Existence of an Analytic Stockwell Transform
The design of an Analytic Stockwell Transform is based on the Cauchy wavelet, also referred to as the Daubechies-Paul wavelet in the literature [12].
@@
Let $\beta\in\mathbb{R}$ be such that $\beta>0$. The Cauchy wavelet of parameters $\beta$ is defined as the unique function $\psi_{\beta}\in L^{2}(\mathbb{R})$ whose Fourier transform reads
@@
Let $\beta>0$ and $\psi_{\beta}\in L^{2}(\mathbb{R})$ be the Cauchy wavelet of parameter $\beta$. Define the modulated Cauchy wavelet $\varphi_{\beta}(t)=\psi_{\beta}(t)\mathrm{e}^{-2\pi\mathrm{i}t}$. There exists a smooth nonvanishing function $\lambda:\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{C}$ and a conformal mapping $\vartheta:\mathbb{C}_{+}\rightarrow\mathbb{D}$ such that any signal $f\in L^{2}(\mathbb{R})$ can be written as
$\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\quad S_{\varphi_{\beta}}f(x,\xi)=\lambda(x,\xi)\times F\left(\vartheta(x+\mathrm{i}\xi^{-1})\right)$ (8)
with $F:\mathbb{D}\rightarrow\mathbb{C}$ an analytic function on the unit disk.
@@
Then, from [27, Section 2.2] and  [21, Theorem 2], the Cauchy Wavelet Transform of $f\in L^{2}(\mathbb{R})$ is analytic on the upper-half complex plane, that is $G:z=x+\mathrm{i}y\mapsto y^{-\beta-1/2}W_{\psi_{\beta}}f(x,y)$ is an analytic function on $\mathbb{C}_{+}$. Moreover, let $\vartheta:\mathbb{C}_{+}\rightarrow\mathbb{D}$ be the Cayley transform defined as
@@
Define $F:\mathbb{D}\rightarrow\mathbb{C}$ as $\forall w\in\mathbb{D},\,F(w)=G(\vartheta^{-1}(w))$. Since $F$ is the composition of an analytic function $G$, and of a conformal, hence holomorphic, mapping $\vartheta^{-1}$, $F$ is analytic. Finally, let $\lambda(x,\xi)=\xi^{-\beta}\mathrm{e}^{-2\pi\mathrm{i}\xi x}$, then $\lambda:\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{C}$ is a smooth nonvanishing function.
By construction of $\vartheta$, $\lambda$ and $F$ they satisfy all the conditions enumerated in Theorem 1 and $\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\,S_{\varphi_{\beta}}f(x,\xi)=\lambda(x,\xi)\times F\left(\vartheta(x+\mathrm{i}\xi^{-1})\right)$ which concludes the proof. ∎
@@ 3 The Zeros of the Analytic Stockwell Transform of White Noise
For the sake of completeness, before proving the connection with the hyperbolic Gaussian analytic function, some necessary elements of hyperbolic and stochastic geometry are presented; the definition of the hyperbolic Gaussian analytic function is reminded accompanied with its main interesting properties so that the present section is self-contained. The interested reader can refer to [22] for a thorough presentation of the theory of Gaussian analytic functions and a detailed study of their zeros.
@@ 3.1 Gaussian analytic functions and point processes
The Poincaré disk model is the triplet $(\mathbb{D},\mathsf{d}_{\mathbb{D}},\mathsf{m}_{\mathbb{D}})$ constituted of the open unit disk $\mathbb{D}=\{z\in\mathbb{C}:\lvert z\rvert<1\}$, the hyperbolic metric defined by
$\mathrm{d}\mathsf{m}_{\mathbb{D}}(z)=\frac{4\mathrm{d}z}{\left(1-\lvert z\rvert^{2}\right)^{2}}$ (12)
where $\mathrm{d}z$ refers to the standard Lebesgue measure on $\mathbb{C}$, and the associated hyperbolic distance defined for $z,w\in\mathbb{D}$ by
@@
Let $\alpha>0$ be a real number, the hyperbolic Gaussian analytic function is defined on the Poincaré disk $\mathbb{D}$ by
@@
where $\left(\zeta_{n}\right)_{n\in\mathbb{N}}$ is a sequence of independent and identically distributed standard Gaussian random variables, $\zeta_{n}\sim\mathcal{N}_{\mathbb{C}}(0,1)$. By [22, Lemma 2.2.3], for $\alpha>0$, $\mathsf{GAF}_{\mathbb{D}}^{(\alpha)}$ is well-defined and almost-surely an analytic function on $\mathbb{D}$.
Gaussian analytic functions have received much interest from the probability and spatial statistics communities, in particular due to the remarkable properties of their zeros [22]. Indeed, since Gaussian analytic functions are almost-surely analytic, with probability one, their zeros constitute a random set of isolated points in $\mathbb{C}$. Further, [22, Lemma 2.4.1] ensures that the random zeros of a Gaussian analytic function are almost-surely simple. Hence, the zero set of a Gaussian analytic function constitutes a simple point process [11, 29], defined as a random variable taking values in the configurations of points in $\mathbb{C}$. Several Gaussian analytic functions, among which the hyperbolic Gaussian analytic function, are distinguished by the fact that their zeros are distributed very uniformly in their definition domain, which is mathematically formalized as an invariance under isometries [22, Section 2.3]. This invariance property is the cornerstone of the zero-based signal processing procedures designed for detection [6, 31], denoising [25, 6] and component separation [16]. The interested reader can refer to [32] for a review of the use of spatial statistics in time–frequency analysis.
@@
[22, Proposition 2.3.4] For $\alpha>0$, the point process formed by the zeros of $\mathsf{GAF}_{\mathbb{D}}^{(\alpha)}$ is invariant under the isometries of the hyperbolic disk.
@@ 3.2 Link with the zeros of the hyperbolic Gaussian analytic function
A major consequence of Theorem 1 is that, since $\vartheta$ is invertible and $\lambda$ never vanishes, the zeros of the Analytic Stockwell Transform of a finite energy signal are in one-to-one correspondance with the zeros of an analytic function $F$ on the unit disk. As so, the zeros of the Analytic Stockwell Transform form a collection of isolated points in $\mathbb{D}$. Further, in the presence of noise in the signal, $f$ is random and then the zeros of $S_{\varphi_{\beta}}f$ are random isolated points in $\mathbb{D}$, and thus form a point process.
In line with [3, 7, 31], the purpose of this section is to study a remarkable instance of such point processes, that is the zeros of the Analytic Stockwell Transform of white noise, and to demonstrate that it coincides in law with the zeros of the hyperbolic Gaussian analytic function introduced in Section 3.1. Following [3, 25, 6, 7], this connection is established in two steps: first the continuous white noise adapted to the considered transform is rigorously constructed; second the equality, in law, of the point process of the zeros of the transform of white noise and of the zeros of the associated Gaussian analytic function is proven. The following lemma will appear as an important building block for the design of the continuous white noise adapted to the Analytic Stockwell Transform.
@@
There exists a smooth nonvanishing function $\eta_{\beta}:\mathbb{R}\times\mathbb{R}_{+}$ such that, for all $n\in\mathbb{N}$, $z=x+\mathrm{i}\xi^{-1}\in\mathbb{C}$, the Analytic Stockwell Transform of $f_{n}$ has the closed-form expression:
@@
The above lemma is an original contribution of the present work. Indeed, first, the orthonormal basis used to connect the Cauchy Wavelet Transform of white noise to the hyperbolic Gaussian analytic function in [3, 25, 7] has to be adapted to the Analytic Stockwell Transform of Section 2.2. Second, in [3, 25] the orthonormal basis used is only defined implicitly. Third, it seems that the explicit derivations performed in [7, Theorem 2.3], relying on [7, Equation (2.15)], together with the computations in [1, Section 4], contains erroneous prefactors that might impact the definition of white noise adapted to the Cauchy Wavelet Transform used to prove [7, Theorem 2.3]. Hence, since Equation (18) is the cornerstone of the demonstration of the connection between the hyperbolic Gaussian analytic function and the Analytic Stockwell Transform of white noise, for the sake of completeness, the proof below details the entire calculation, including the proposed corrected prefactors and normalizations.
@@
involves the convolution between $f$ and $h$ defined by $h(t)=\overline{\psi_{\beta}(-t/\xi^{-1})}$. Then, using that the Fourier transform of the convolution between two functions is the product of the Fourier transforms of these functions leads to
@@
which is the Laplace transform [^1] Let $g:\mathbb{R}\rightarrow\mathbb{C}$. The Laplace transform of $g$, denoted $\mathcal{L}\{g\}:\mu\subseteq\mathbb{C}\rightarrow\mathbb{C}$, is a function of the complex variable $s$ defined by: $\mathcal{L}\{g\}(s)=\int_{\mathbb{R}}\mathrm{e}^{-\nu s}g(\nu)\,\mathrm{d}\nu$, $s\in\mu$. of the $n$th order derivative of $g_{n}(\nu)=\mathrm{e}^{-\nu}\nu^{2\beta+n}$ evaluated at $-(\mathrm{i}z+1)/2$, i.e.,
@@
$\begin{split}\mathcal{L}\{g_{n}\}(s)=\int_{0}^{\infty}\mathrm{e}^{-s\nu}\mathrm{e}^{-\nu}\nu^{2\beta+n}\,\mathrm{d}\nu&=\frac{\Gamma(2\beta+n+1)}{(1+s)^{2\beta+n+1}}\end{split}$ (25)
leveraging the change of variable $\nu^{\prime}=\nu(1+s)$ and the definition of the gamma function. The second step is to compute the limits $g^{(k)}(0^{+})$ for $k\in\{0,\ldots,n-1\}$. To that aim, a recursion shows that for all $k\in\{0,1,\ldots,n\}$, there exists a polynomial $p_{k}$ of order $n$, such that
@@
where $\vartheta$ is the Cayley transform defined in Equation (10). Finally, remembering that $z=x+\mathrm{i}\xi^{-1}$, let $\eta_{\beta}:\mathbb{R}\times\mathbb{R}_{+}\rightarrow\mathbb{C}$ be defined by
$\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\quad\eta_{\beta}(x,\xi)=\sqrt{\xi}\mathrm{e}^{-2\pi\mathrm{i}\xi x}\left(\frac{\mathrm{i}}{x+\mathrm{i}\xi^{-1}+\mathrm{i}}\right)^{2\beta+1}.$ (29)
Then, for any $\beta>0$, $\eta_{\beta}$ is a well-defined smooth nonvanishing function on the upper half-plane, and the Analytic Stockwell Transform of $f_{n}$ writes